A river 4 miles wide is following at the rate of 2 miles /hr. The minimum time taken by a boat to cross the river with a speed of 4 miles/hr is approximately

Grade:10

1 Answers

John Sharon Sandeep
6 years ago
Here speed of river=2mph. Speed of boat in still water=4mph. Therefore speed of the boat in the flowing river=2mph+4mph=6mph. Since distance=4miles , by S=D/T , 6mph=4m/T => T= 4/6 h = 40 minutes.☺